UT 20150401 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, WSeifert, APramskiy, DThompson (LBTO)
Observer AO-Team: SEsposito, GAgapito, EPinna, APuglisi
Telescope Operator: GBechetti
LBTO Support: TSargent

Plan for the night

LUCI/AgW still in the lab => This night has been converted to a LBTO engineering night

Summary

LBTO engineering night was good

Did some detector tests and focus series with LUCI2 in the lab

Arcetri group made progress with their AGw/AO problem

G40 flexure issue, found workaround

Preparations

Details

All times are UT
Files are luci2.20150401.NNNN.fits

  • Focus LUCI2

05:16 N3.75, K, spectro sieve, start = -350, step = 5, Nsteps = 11, #2-12, 3sec wrong start/step (all out of focus)
redo, start= -450, step = 10, Nsteps = 11, #13-23, 3sec, OK

N1.8, K, spectro sieve start = -350, step = 5, Nsteps = 11, 3sec, #24-34

N30, K, N30 sieb II, start = -500, step = 50, Nsteps = 11, 30sec, #35-45

  • Detector tests

Idle-wait, type-lir

N3.75, Ks, DIT=3sec, NDIT=1, NExp=20, #46-65

N3.75, Ks, DIT=3sec, NDIT=20, NExp=1, #66-85

Test for overheads:

Idle-wait, type-lir, DIT=20, NExp=20, #86-105, dT = 627sec

Idle-break, type=rwc, DIT=20, NExp=20, #106-125, dT = 605sec

Idle-wait, type=rwc, DIT=20, NExp=20, #126-145, dT = 847sec

=> following RMathars proposal using WAIT type LIR
-> step function not present, total time ~ break
-> success!!!!

G40 flexure issue
apparently the notch is not working correctly
shifts of up to several hundreds pixels observed
only moving from G200 by +-249steps ok
(before +252/-253steps)
How to do: move to G40, then G200, then G40
Topic revision: r8 - 02 Apr 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback